OpenStudy (anonymous):

Triangle ABC has the following side lengths: AB = 5 ft, BC = 10 ft, CA = 12 ft Triangle DEF is similar to triangle ABC. The shortest side of triangle DEF, DE = 8.5 ft. What are the lengths for the other two sides of triangle DEF? A. EF 17 ft; FD = 20.4 ft B. EF = 6 ft; FD = 7.2 ft C. EF = 10 ft; FD = 12 ft D. EF = 20 ft; FD = 24 ft

OpenStudy (anonymous):

Set up a proportion. Find the corresponding sides.

Directrix (directrix):

Triangle DEF is similar to triangle ABC. Lengths of corresponding sides of similar triangles are in proportion. Given the triangle similarity statement, AB/DE = AC/DF = BC/EF Using the given values, 5/8.5 = 12/DF = 10/EF Take the first two equal ratios, 5/8.5 = 12/DF Cross multiply and solve for DF DF = 20.4 Then, do the same for the equal ratios to solve for EF. 5/8.5 =10/EF @arlira